OpenStudy (ronlover101):

Please help I will medal What is the slope of the line? y=2x-7 slope =

OpenStudy (anonymous):

2x

OpenStudy (ronlover101):

Can you go throw the steps with me?

OpenStudy (ronlover101):

Cause I don't know how to do that.

OpenStudy (anonymous):

There really isn't any steps.......x is your slope sooooooo.....in this case.....its 2x

OpenStudy (anonymous):

if there's is a reciprical(number)......always include it

OpenStudy (ronlover101):

I don't know if that is right.

OpenStudy (ronlover101):

but ok

OpenStudy (anonymous):

XD......nvm.......its just 2

OpenStudy (anonymous):

lol.......looking the wrong way......

OpenStudy (anonymous):

https://mathway.com/

OpenStudy (ronlover101):

What is the y-intercept of the line? y=-2x-8 y-intercept =

OpenStudy (anonymous):

Great site to refer to btw

OpenStudy (ronlover101):

Thanks

OpenStudy (anonymous):

y-intercept: (0,−8)

OpenStudy (ronlover101):

What is the slope, and what is the y-intercept of the line? y=2/3x-4 slope = ?/? y-intercept =

OpenStudy (anonymous):

2/3=slope y-intercept: (0,−4)

OpenStudy (ronlover101):

thanks

OpenStudy (ronlover101):

Question Navigator 4. Transform this equation to slope-intercept form: 6x + 4y = 8. A. y equals negative begin fraction three over two end fraction times x plus four B. y equals negative begin fraction three over two end fraction times x plus two C. y equals begin fraction three over two end fraction times x plus four D. y equals begin fraction three over four end fraction times x plus two Question Resources

OpenStudy (anonymous):

yeah!I'll fan you so that if you need any questions and you want me to answerit you can ask lol alott of ppl are doin that to me

OpenStudy (ronlover101):

Can anyone answer my question.

OpenStudy (anonymous):

y = 2x + 1 and y = -x + 7. 1. The equations in this example are already set equal to "y". Identify the slope and y-intercept in each equation.

OpenStudy (ronlover101):

so (2,1) for line 2, and (2, -1) for line 1 is the answer?

OpenStudy (anonymous):

yes

OpenStudy (ronlover101):

Ok thanks

OpenStudy (anonymous):

welcome <3

OpenStudy (ronlover101):

@CheesecakeKitten why did you post those on my question. Than is VERY RUDE!!!!

OpenStudy (anonymous):

i agree with @ronlover101